# Capturing Human Motion from Monocular Images in World Space with Weak-supervised Calibration

### v4

@@ III-B Weak-Supervised Camera Calibration
Weak Supervision Our main losses can be categorized into joint loss, vertex loss, and parameter loss. In this section, we only introduce the joint loss. Please refer to Sec. III-E for information on the others. The joint loss can be calculated using the following function:
@@
where $MSE(\cdot)$ represents mean square error. The $\land$ here denotes ground truth. The $J_{3D}$ denotes 3D joints, which are regressed from mesh vertices[1]. $J_{2D}$ signifies normalized 2D joint coordinate in the cropped image and ${J}_{2D}^{full}$ is normalized 2D joint coordinate in the full image. Note that 2D joints are the projected 3D joints on images. $J_{2D}$ and $J_{2D}^{full}$ are the same things but based on different coordinates (cropped image vs full image). The last term in the Equ. 5 is the key of our FULL2 2D supervision. It is worth mentioning that $\left(\frac{W}{w_{bbox}},\frac{H}{h_{bbox}}\right)$ is used to dynamically balance the weights and prevent the FULL2 2D loss from being too small, which could lead to ineffective supervision.
Figure 5: This figure depicts our model training paradigm, illustrating the detachment strategy at different stages to achieve stable model training. The symbol $\otimes$ represents detachment. Traditional 2D supervision refers to projecting 3D joints onto the cropped image to get 2D joints based on the weak-perspective camera model, just as traditional models did. The ”freeze” means other modules are frozen in the third stage, and only OrientCorrect is updated.
Weak supervision is reflected in this FULL2 2D supervision. We do not use any real focal length label to supervise the training of the focal length predictor. From Sec. III-A, we know that the focal length is involved in the 3D to 2D projection. So, by having 2D and 3D joint labels, we can let the predictor learn how to generate a correct focal length for 3D to 2D joint projection. After all, all this effort to make the projection more precise is so that the model can be trained more accurately. By the way, rich joint labels are easier to collect for model training than rare focal length labels.
III-C Orientation Correction
@@
where $\oplus$ represents concatenation, $FC(\cdot)$ are only couples of FC layers, $R_{c}$ is the camera rotation output by CamClib, $R_{b}^{c}$ is the predicted body orientation $\theta_{0}$, and $R_{b}$ is the final predicted body orientation in the world coordinate. Under the supervision of true body orientation in the world coordinate, $FC(\cdot)$ learns to correct the $R_{b}^{c}$ to $R_{b}$. No additional operation is required to ensure the output rotation matrix $R_{b}$ is valid. To avoid distress, we explain why the operation of rotation matrices here does not use multiplication but rather direct addition. The network of OrientCorrect is similar to the classic iterative regressor used in HMR [1]. A similar design, instead of multiplication, has been employed in many previous works [4, 11, 13, 34] to predict rotations. The global information $glob\_info=\left\{\frac{\varDelta c_{x}}{\tilde{f}},\frac{\varDelta c_{y}}{\tilde{f}},\frac{h_{bbox}}{\tilde{f}},\frac{W}{\tilde{f}},\frac{H}{\tilde{f}}\,\,\right\}$ and it is important auxiliary information for predicting body orientation [4]. As for $\tilde{f}=\sqrt{W^{2}+H^{2}}$ in global information, it is just for normalization here and does not hold any geometric meanings, so we do not use the predicted focal length here. The OrientCorrect is a completely separate module. So, it not only makes full use of what the model has learned in the camera coordinate but also avoids destroying the body feature space that the model has constructed. Additionally, this method can be easily applied to other models based on the camera coordinate without reducing the original precision, shown in Tab. VIII.
III-D Training Paradigm
@@ III-E Other Losses
To enhance the model’s ability to focus on human regions and extract refined body information, we include two additional sub-tasks for auxiliary supervision, which are inspired by [13].
First, let’s delve into the IUV map, which is integral to the IUV loss. The IUV map, as defined in [35], uses ”I” to indicate the body part to which a pixel belongs and ”UV” to represent the 2D coordinates on the surface of the body part. This mapping establishes dense correspondences from 3D mesh to 2D plane. We employ three 2D convolution layers to predict the IUV map based on the feature map $\varPhi_{3}$ and render the ground truth online. The loss function is defined as
@@
TABLE I: This table illustrates the changes in the size of input images and features when using different backbones (ResNet50 or VitPose-B) and employing different downsampling methods (sparse vertices or markers). As for what symbols mean, please refer to Sec. IV-A.
ResNet50 + vertices ResNet50 + markers VitPose-B + vertices VitPose-B + markers
$H_{c}\times W_{c}$ 224$\times$224 224$\times$224 256$\times$192 256$\times$192
[…]
$C_{i}^{fc}$ 5 32 5 32
$H_{g}\times W_{g}$ 21$\times$21 8$\times$8 24$\times$18 9$\times$7
$\{C_{i}^{s}\}_{i=1}^{3}$2205,2155,2155} {2048,2144,2144} {2160,2155,2155} {2016,2144,2144}
IV Experiment
IV-A Implementation Details
Pixel-aligned Feature Extractor We follow the feature extraction method of [13] but use different sampling points and backbones. When a cropped image $I\in\mathbb{R}^{H_{c}\times W_{c}}$ is inputted into the encoder, it generates an initial feature map $\varPhi_{0}\in\mathbb{R}^{C\times H_{0}\times W_{0}}$. Subsequently, this feature map $\varPhi_{0}$ undergoes further processing through a series of deconvolution layers to yield high-resolution features $\left\{\varPhi_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}\right\}_{i=1}^{3}$. For the first high-resolution feature map $\varPhi_{1}$, a $H_{g}\times W_{g}$ grid is defined to sample finer point-wise features. But, for the second and third feature maps, we project vertices of the mesh obtained from the previous step onto the feature map to sample point-wise features. In summary, we extract spatial features based on the following formula.
[…]
where $P_{i}$ represents the projected 2D point for sampling, $bs(\cdot)$ denotes bilinear sampling, and $fc(\cdot)$ refers to FC layers for reducing channel size. Finally, $\oplus$ signifies the concatenation of individual point-wise features $\left\{\phi_{i,n}^{p,fc}\in\mathbb{R}^{C_{i}^{fc}}\right\}_{n=1}^{N_{i}}$ into a one-dimensional feature $\phi_{i}^{s}\in\mathbb{R}^{C_{i}^{s}}$, where $C_{i}^{s}=N_{i}\times C_{i}^{fc}$. $fc$ indicates the point-wise features has been through $fc(\cdot)$ for channel reduction and $N_{i}$ is the number of points for downsampling.
Hybrid Regression Module For the spatial features $\{\phi_{i}^{s}\}_{i=1}^{3}$, we have corresponding regression models. We utilize parameter regressors for the first two spatial features to obtain the initial motion. These regressors take spatial features $\phi_{i}^{s}$, global information $glob\_info$, and former SMPL parameters $\varTheta_{i-1}$ as inputs and output SMPL parameters $\varTheta_{i}$. The parameter regressors consist of several FC layers and iteratively produce the final result like
$\varTheta_{i}=\varTheta_{i-1}+FC\left(\oplus\left(\phi_{i}^{s},\,\,glob\_info,\ \varTheta_{i-1}\right)\right)$ (13)
Note $\varTheta_{0}$ is set as the mean $\overline{\varTheta}$ following [1]. As for the final regressor, we have two versions in total. The first version still involves the parameter regressor. The second version utilizes a vertex regressor to obtain a finer human body motion. After obtaining a coarse human motion using the first two regressors, we introduce a model-free vertex regressor based on the transformer. This vertex regressor takes spatial features $\phi_{3}^{s}$ and point-wise features $\left\{\phi_{3,n}^{p}\in\mathbb{R}^{C_{i}}\right\}_{n=1}^{N_{i}}$ as input. It’s important to note that no ”$fc$” exists in this case, indicating that these features do not undergo channel reduction. To reduce network complexity and adapt to the SMPL model, instead of directly recovering the 6890 vertices defined by SMPL, we recover 431 sparse vertices and gradually restore them to 6890 vertices using two linear layers.
@@ IV-A Implementation Details
We train three models in total for evaluation. Two of them utilize a hybrid regression module, using 431 sparse vertices for downsampling. The distinction between these two models lies in using different backbones, including ResNet50 [36] and VitPose-B [37]. Additionally, we train an extra model using a pure parameter regression module, employing 67 mocap markers for downsampling, with the backbone being VitPose-B. For training, we utilize datasets such as Human 3.6m [38], 3DPW [39], MPII3D [40], MPII [41], AGORA [42], COCO [43], SPEC-SYN [3], and HuMMan [44]. Some datasets are augmented with pseudo-labels generated by CamCalib, EFT [45], or CLIFF [4]. Please refer to Sec. IV-B for detailed information about the datasets used. Our model can simultaneously provide results in camera and world coordinates. But if results in the world coordinate are not needed or camera rotation is known, the CamCalib and OrientCorrect can be discarded. The focal length predictor can also be discarded if mesh projection is not required.
We experiment with different backbones and downsampling points to validate our method and explore the optimal model structure. As a result, there are corresponding variations in the sizes of input images and features. We summarize these variations and create Tab. I to provide an overview of these changes. The training process is conducted with a batch size of 128 and a fixed learning rate of 0.00005 on three RTX 3090 GPUs until convergence, taking approximately one week. Following [13, 11], we apply standard data augmentation techniques such as rotation, flipping, scaling, translation, noise, and occlusion. As for hyperparameters, $\lambda_{1}=\lambda_{2}=\lambda_{3}=300$, $\lambda_{4}=\lambda_{5}=\lambda_{6}=15$, $\lambda_{7}=0.06$, $\lambda_{8}=60$, $\lambda_{9}=0.03$ and $\lambda_{10}=\lambda_{11}=\lambda_{12}=0.05$.
IV-B About Datasets
In our experiments, we utilize the following datasets: Human 3.6M [38], 3DPW [39], MPII [41], COCO [43], MPI-INF-3D [40], AVA  [46], AIC [47], Insta [48], AGORA [42], SPEC-SYN [3], SPEC-MTP [3], and HuMMan [44]. To evaluate our model, we conduct experiments on the distorted datasets AGORA, HuMMan, and SPEC-MTP. Furthermore, following traditional practices[49, 13], we also evaluate our model on 3DPW, Human 3.6M, and MPII-INF-3D. We employ AGORA’s evaluation platform as a standard for ablation experiments due to its diverse scenes, characters, and shooting methods. We believe that the evaluation results obtained on AGORA are more representative. For human recovery in world space, We additionally perform an ablation study on SPEC-SYN to validate the robustness of our method. Following are details about each dataset.
3DPW’ [39] full name is ”3D Human in the Wild.” It is an outdoor dataset that captures actor information using motion capture devices. This dataset is known for its difficulty and has been widely used by researchers to demonstrate the robustness of their methods. Following past conventions, we train two models, one without 3DPW for training and another using 3DPW for training, to evaluate our model’s robustness.
Human 3.6M [38] is widely used as a standard dataset to assess the estimation of 3D human pose and the recovery of human motion. To reduce data redundancy, the original videos are down-sampled from 50 to 10 fps, resulting in 312,188 frames available for training and 26859 frames for evaluation. Following established protocols, we utilize five subjects (S1, S5, S6, S7, S8) for training and two subjects (S9, S11) for evaluation.
MPII [41] is collected from YouTube and comprises various activities in various scenes. Based on [7, 14], we only utilize the portion with pseudo 3D labels, containing approximately 14,000 images.
COCO [43] is a large-scale image dataset widely used for various computer vision tasks such as object detection, image segmentation, and image classification. For our task, we primarily utilize the part including human subjects. COCO provides 2D joint location labels. Based on this, we use corresponding pseudo 3D labels generated by [4, 45, 50], following [7, 4, 14].
MPI-INF-3D [40, 51] is a large-scale 3D dataset that includes both indoor and outdoor scenes. To reduce data redundancy, we extract every 10th frame. The dataset is divided into a training set and a test set according to their original partitioning. The training set comprises 96,507 images, while the test set contains 2929 images. We use pseudo SMPL parameter labels generated by [45].
@@ IV-B About Datasets
TABLE II: Comparison with SOTA methods on distorted datasets in the camera coordinate. Some of the evaluation results on HuMMan [44] and SPEC-MTP [3] are sourced from  [2], while the evaluation results on AGORA [42] are obtained from its official website https://agora-evaluation.is.tuebingen.mpg.de/. Including the following tables, represents backbone is ResNet50 [36], means VitPose-B [37, 52], and ”P” denotes the pure parameter module. Note we do not perform any fine-tuning on SPEC-MTP.
@@
HMR [1] 217.0 226.0 173.6 180.5 43.6 30.2 52.6 121.4 73.9 145.6
HMR-f [1] - - - - 43.6 29.9 53.4 123.2 72.7 145.1
SPEC [3] 126.8 133.7 106.5 112.3 44.0 31.4 54.2 125.5 76.0 144.6
CLIFF [4] 83.5 89.0 76.0 81.0 42.4 28.6 50.2 115.0 74.3 132.4
PARE [11] 167.7 174.0 140.9 146.2 53.2 32.6 65.5 121.6 74.2 143.6
GraphCMR [53] - - - - 40.6 29.5 48.4 121.4 76.1 141.6
FastMETRO [54] - - - - 38.8 26.3 45.5 123.1 75.0 137.0
PLIKS [29] 71.6 76.1 67.3 71.5 - - - - - -
ProPose [55] 78.8 82.7 70.9 74.4 - - - - - -
HybrIK [56] 81.2 84.6 73.9 77.0 - - - - - -
Hand4Whole [57] 90.2 95.5 84.8 89.8 - - - - - -
PyMAF [13] 87.3 92.6 78.6 83.3 42.5 28.9 50.4 108.6 66.1 127.5
Zolly [2] - - - - 32.6 22.3 40.0 114.6 67.4 126.7
@@
Figure 6: Qualitative results on AGORA compared with SPEC [3]. The areas marked within the red circles indicate the mismatch between the SPEC reconstructed motion and the original image. In contrast, by comparing it with the corresponding W-HMR reconstructed results enclosed in the yellow circles, it is evident that our approach exhibits stronger capabilities in handling complex samples.
AVA [46], AI Challenger[47] and InstaVariety [48] are three huge wild datasets that lack real 3D labels. However, in the Human4D [14], researchers utilized VitPose [37] to obtain 2D keypoints and VitDet [58] to generate bounding boxes for human targets. Based on these 2D pseudo labels, they employed ProHMR [50] to acquire pseudo SMPL labels. Subsequently, we further optimize these pseudo-labeled datasets by selecting high-quality samples for training our model.
AGORA [42] is a multi-person synthetic dataset that has gained recognition recently due to its precise SMPL-X labels, complex characters and scenes, and fair evaluation methods. The dataset is available in two resolutions: 1280x720 and 3840x2160. We exclusively utilize the low-resolution set in our work. The train, validation, and test sets are used according to their original partitioning. The bounding boxes used in the test set are obtained using the 2D keypoints generated by Hand4Whole [57].
SPEC-SYN and SPEC-MTP are datasets proposed in SPEC [3]. Inspired by AGORA, SPEC-SYN is also a synthetic dataset with 3,783 images synthesized from five large-scale virtual scenes. It offers a relatively high diversity of focal lengths. On the other hand, SPEC-MTP is a real dataset with SMPL-X labels generated by multi-view fitting. It consists of 3,284 images, which we solely use for evaluating our model. The dataset is challenging because of its serious distortion and special shooting angle.
HuMMan[44] is a high-precision human dataset constructed indoors. It includes labels such as keypoints, point clouds, meshes, SMPL parameters, textures, and more. We only utilize its SMPL labels. We sample every five images to form a training set comprising approximately 84,000 images. Due to the short focal length during data collection, it contains a significant amount of distorted data. Therefore, we also use HuMMan as one of our evaluation datasets. The partitioning of the train and test sets follows the original division provided by the dataset creators.
IV-C About Metrics
@@
We evaluate the reconstruction accuracy using 3D mesh vertices and joints. Specifically, for AGORA, we employ four metrics: MPJPE, MVE, NMJE, and NMVE. MPJPE stands for Mean Per Joint Position Error, which measures the accuracy of joint positions. For AGORA, it calculates the error for the first 24 SMPL joints after aligning with the root joint. Following [13, 9], we utilize 17 joints for MPI-INF-3D, while the other datasets employ a consistent set of 14 keypoints. MVE, or Mean Vertex Error, measures the error of all SMPL mesh vertices after aligning the root vertices. PVE (Per Vertex Error) is computed without alignment for other datasets except AGORA. These metrics help assess the accuracy of the reconstructed human bodies. NMJE (Normalized Mean Joint Error) and NMVE (Normalized Mean Vertex Error) are obtained by normalizing MPJPE and MVE using the F1 score. They provide an overall evaluation of the method’s performance, particularly in handling challenging scenarios. For specific calculation methods, please refer to [42].
@@
As for the metrics in the world coordinate, we follow SPEC [3] using W-MPJPE, PA-MPJPE, and W-PVE. PA-MPJPE (Procrustes-aligned Mean Per Joint Position Error) exists as a metric that purely evaluates the accuracy of body reconstruction. It removes the effect of unknown camera rotation by Procrustes-aligned and has therefore been adopted by many previous works. The calculation of W-MPJPE and W-PVE is the same as that of MPJPE and PVE. The only difference is that the labels used are adjusted to the ground truth in the world coordinate. We, including SPEC, believe these metrics represent the model’s capabilities in real-world applications.
IV-D Evaluation Results

### v6

@@ 3.2 Weak-Supervised Camera Calibration
Weak Supervision Our main losses can be categorized into joint loss, vertex loss, and parameter loss. In this section, we only introduce the joint loss. Please refer to Sec. 3.5 for information on the others. The joint loss can be calculated using the following function:
@@
where $MSE(\cdot)$ represents mean square error. The $\land$ here denotes ground truth. The $J_{3D}$ denotes 3D joints, which are regressed from mesh vertices(Kanazawa et al., 2018). $J_{2D}$ signifies normalized 2D joint coordinate in the cropped image and ${J}_{2D}^{full}$ is normalized 2D joint coordinate in the full image. Note that 2D joints are the projected 3D joints on images. $J_{2D}$ and $J_{2D}^{full}$ are the same things but based on different coordinates (cropped image vs full image). The last term in the Equ. 5 is the key of our FULL2 2D supervision. It is worth mentioning that $\left(\frac{W}{w_{bbox}},\frac{H}{h_{bbox}}\right)$ is used to dynamically balance the weights and prevent the FULL2 2D loss from being too small, which could lead to ineffective supervision.
Weak supervision is reflected in this FULL2 2D supervision. We do not use any real focal length label to supervise the training of the focal length predictor. From Sec. 3.1, we know that the focal length is involved in the 3D to 2D projection. So, by having 2D and 3D joint labels, we can let the predictor learn how to generate a correct focal length for 3D to 2D joint projection. After all, all this effort to make the projection more precise is so that the model can be trained more accurately. By the way, rich joint labels are easier to collect for model training than rare focal length labels.
3.3 Orientation Correction
@@
where $\oplus$ represents concatenation, $FC(\cdot)$ are only couples of FC layers, $R_{c}$ is the camera rotation output by CamClib, $R_{b}^{c}$ is the predicted body orientation $\theta_{0}$, and $R_{b}$ is the final predicted body orientation in the world coordinate. Under the supervision of true body orientation in the world coordinate, $FC(\cdot)$ learns to correct the $R_{b}^{c}$ to $R_{b}$. No additional operation is required to ensure the output rotation matrix $R_{b}$ is valid. To avoid distress, we explain why the operation of rotation matrices here does not use multiplication but rather direct addition. The network of OrientCorrect is similar to the classic iterative regressor used in HMR (Kanazawa et al., 2018). A similar design, instead of multiplication, has been employed in many previous works (Li et al., 2022b; Kocabas et al., 2021a; Zhang et al., 2021a; Wang and Daniilidis, 2023) to predict rotations. The global information
$glob\_info=\left\{\frac{\varDelta c_{x}}{\tilde{f}},\frac{\varDelta c_{y}}{\tilde{f}},\frac{h_{bbox}}{\tilde{f}},\frac{W}{\tilde{f}},\frac{H}{\tilde{f}}\,\,\right\}$ (7)
is important auxiliary information for predicting body orientation (Li et al., 2022b). As for $\tilde{f}=\sqrt{W^{2}+H^{2}}$ in global information, it is just for normalization here and does not hold any geometric meanings, so we do not use the predicted focal length here. The OrientCorrect is a completely separate module. So, it not only makes full use of what the model has learned in the camera coordinate but also avoids destroying the body feature space that the model has constructed. Additionally, this method can be easily applied to other models based on the camera coordinate without reducing the original precision, shown in Tab. 7.
Figure 5: This figure depicts our model training paradigm, illustrating the detachment strategy at different stages to achieve stable model training. The symbol $\otimes$ represents detachment. Traditional 2D supervision refers to projecting 3D joints onto the cropped image to get 2D joints based on the weak-perspective camera model, just as traditional models did. The “freeze” means other modules are frozen in the third stage, and only OrientCorrect is updated.
3.4 Training Paradigm
@@ 3.5 Other Losses
To enhance the model’s ability to focus on human regions and extract refined body information, we include two additional sub-tasks for auxiliary supervision, which are inspired by (Zhang et al., 2021a).
First, let’s delve into the IUV map, which is integral to the IUV loss. The IUV map, as defined in (Güler et al., 2018), uses “I” to indicate the body part to which a pixel belongs and “UV” to represent the 2D coordinates on the surface of the body part. This mapping establishes dense correspondences from 3D mesh to 2D plane. We employ three 2D convolution layers to predict the IUV map based on the feature map $\varPhi_{3}$ and render the ground truth online. The loss function is defined as
@@
4 Experiment
4.1 Implementation Details
@@
We train three models in total for evaluation. Two of them utilize a hybrid regression module, using 431 sparse vertices for downsampling. The distinction between these two models lies in using different backbones, including ResNet50 (He et al., 2016) and VitPose-B (Xu et al., 2022a). Additionally, we train an extra model using a pure parameter regression module, employing 67 mocap markers for downsampling, with the backbone being VitPose-B. For training, we utilize datasets such as Human 3.6m (Ionescu et al., 2013), 3DPW (Von Marcard et al., 2018), MPII3D (Mehta et al., 2017a), MPII (Andriluka et al., 2014), AGORA (Patel et al., 2021), COCO (Lin et al., 2014), SPEC-SYN (Kocabas et al., 2021b), and HuMMan (Cai et al., 2022). Some datasets are augmented with pseudo-labels generated by CamCalib, EFT (Joo et al., 2021), or CLIFF (Li et al., 2022b). Please refer to appendices for detailed information about the datasets used. Our model can simultaneously provide results in camera and world coordinates. But if results in the world coordinate are not needed or camera rotation is known, the CamCalib and OrientCorrect can be discarded. The focal length predictor can also be discarded if mesh projection is not required.
@@
Table 1: Comparison with SOTA methods on distorted datasets in the camera coordinate. Some of the evaluation results on HuMMan (Cai et al., 2022) and SPEC-MTP (Kocabas et al., 2021b) are sourced from  (Wang et al., 2023), while the evaluation results on AGORA (Patel et al., 2021) are obtained from its official website https://agora-evaluation.is.tuebingen.mpg.de/. Including the following tables, represents backbone is ResNet50 (He et al., 2016), means VitPose-B (Xu et al., 2022a, b), and “P” denotes the pure parameter module. Note we do not perform any fine-tuning on SPEC-MTP.
@@
HMR (Kanazawa et al., 2018) 217.0 226.0 173.6 180.5 43.6 30.2 52.6 121.4 73.9 145.6
HMR-f (Kanazawa et al., 2018) - - - - 43.6 29.9 53.4 123.2 72.7 145.1
SPEC (Kocabas et al., 2021b) 126.8 133.7 106.5 112.3 44.0 31.4 54.2 125.5 76.0 144.6
CLIFF (Li et al., 2022b) 83.5 89.0 76.0 81.0 42.4 28.6 50.2 115.0 74.3 132.4
PARE (Kocabas et al., 2021a) 167.7 174.0 140.9 146.2 53.2 32.6 65.5 121.6 74.2 143.6
GraphCMR (Kolotouros et al., 2019b) - - - - 40.6 29.5 48.4 121.4 76.1 141.6
FastMETRO (Cho et al., 2022) - - - - 38.8 26.3 45.5 123.1 75.0 137.0
PLIKS (Shetty et al., 2023) 71.6 76.1 67.3 71.5 - - - - - -
ProPose (Fang et al., 2023) 78.8 82.7 70.9 74.4 - - - - - -
HybrIK (Li et al., 2021) 81.2 84.6 73.9 77.0 - - - - - -
Hand4Whole (Moon et al., 2022) 90.2 95.5 84.8 89.8 - - - - - -
PyMAF (Zhang et al., 2021a) 87.3 92.6 78.6 83.3 42.5 28.9 50.4 108.6 66.1 127.5
Zolly (Wang et al., 2023) - - - - 32.6 22.3 40.0 114.6 67.4 126.7
@@
We experiment with different backbones and downsampling points to validate our method and explore the optimal model structure. As a result, there are corresponding variations in the sizes of input images and features. We summarize these variations in the appendices to provide an overview of these changes. The training process is conducted with a batch size of 128 and a fixed learning rate of 0.00005 on three RTX 3090 GPUs until convergence, taking approximately one week. Following (Zhang et al., 2021a; Kocabas et al., 2021a), we apply standard data augmentation techniques such as rotation, flipping, scaling, translation, noise, and occlusion. As for hyperparameters, $\lambda_{1}=\lambda_{2}=\lambda_{3}=300$, $\lambda_{4}=\lambda_{5}=\lambda_{6}=15$, $\lambda_{7}=0.06$, $\lambda_{8}=60$, $\lambda_{9}=0.03$ and $\lambda_{10}=\lambda_{11}=\lambda_{12}=0.05$.
4.2 About Metrics
@@
We evaluate the reconstruction accuracy using 3D mesh vertices and joints. Specifically, for AGORA, we employ four metrics: MPJPE, MVE, NMJE, and NMVE. MPJPE stands for Mean Per Joint Position Error, which measures the accuracy of joint positions. For AGORA, it calculates the error for the first 24 SMPL joints after aligning with the root joint. Following (Zhang et al., 2021a; Kolotouros et al., 2019a), we utilize 17 joints for MPI-INF-3D, while the other datasets employ a consistent set of 14 keypoints. MVE, or Mean Vertex Error, measures the error of all SMPL mesh vertices after aligning the root vertices. PVE (Per Vertex Error) is computed without alignment for other datasets except AGORA. These metrics help assess the accuracy of the reconstructed human bodies. NMJE (Normalized Mean Joint Error) and NMVE (Normalized Mean Vertex Error) are obtained by normalizing MPJPE and MVE using the F1 score. They provide an overall evaluation of the method’s performance, particularly in handling challenging scenarios. For specific calculation methods, please refer to (Patel et al., 2021).
@@
As for the metrics in the world coordinate, we follow SPEC (Kocabas et al., 2021b) using W-MPJPE, PA-MPJPE, and W-PVE. PA-MPJPE (Procrustes-aligned Mean Per Joint Position Error) exists as a metric that purely evaluates the accuracy of body reconstruction. It removes the effect of unknown camera rotation by Procrustes-aligned and has therefore been adopted by many previous works. The calculation of W-MPJPE and W-PVE is the same as that of MPJPE and PVE. The only difference is that the labels used are adjusted to the ground truth in the world coordinate. We, including SPEC, believe these metrics represent the model’s capabilities in real-world applications.
4.3 Evaluation Results